OpenStudy (anonymous):

Use the graph below to fill in the blank with the correct number. f(1) = _______

OpenStudy (anonymous):

Here is the graph

OpenStudy (anonymous):

@HelpMePlease?

OpenStudy (anonymous):

Are there choices?

OpenStudy (anonymous):

no I have to fill in the blank

OpenStudy (anonymous):

can u explain how u got -2?

OpenStudy (anonymous):

what do you think it is @helpme1.2

OpenStudy (anonymous):

i dont really understand the question the points are (1,-2)(0,1)(-3,0)and (-2.2) what is the bank for

OpenStudy (anonymous):

i have to fill in a number from the graph that is f(1)

OpenStudy (anonymous):

ooh f(1) is defined as f(x), if so , then cookie_scott is correct, it's -2

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

Is the answer for this 8 Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (anonymous):

f(1) = 2 and f(2) = 5 is the given solution And f(n) = f(1) + f(2) + f(n - 1), for n > 2. is the given form?

OpenStudy (anonymous):

f(5)=f(5)+f(5)+f(5-1) = the answer

OpenStudy (anonymous):

which would be 14